• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Sirca

    Log entry time 00:00:59 on March22,2005

    Entry number 142142

    keyword=Swing shift summary

    SWING Shift Crew: S. Sirca (SL), B. Anderson (TO)


    EVENTS

    16:00 Can not get beam within desired position/spread tolerances
    18:25 Hall access to reset NDet top HV crate
    19:00 Hall back to Beam Permit
    19:05 Beam back, attempting to optimize (position, width) and make harp scans
    20:25 Beam back, resume production running
    22:40 Beamline vacuum problems, reducing current to 20 uA temporarily, see 142138
    23:05 Increase I_beam to 30 uA, see 142139
    23:20 Hall A laser lost lock, restored in 5 minutes, called injector-on-call


    RUN HISTORY

    #2855 ..... junk
    #2856 ..... junk
    #2857 ..... junk
    #2858 ..... 200K
    #2859 ..... 200K
    #2860 ..... 140K
    #2861 ..... 80K (run with 20 uA in first 50K events, rest with 30 uA)
    #2862 ..... passed on to the nightshift


    COMMENTS

    Attempts by the MCC from ~19h to ~20h to steer the beam properly knocked out
    a random pattern of HV channels in several slots of different crates; we were
    lucky enough to be able to restore them all remotely.